OpenStudy (anonymous):

Find the slope of the line that passes through the points (2, -3) and (-5,1). a. -4/7 b. -2/3 c. 1/2 d. 3/7

OpenStudy (anonymous):

a

OpenStudy (anonymous):

Slope is easy, and it's fun. Just remember rise over run!

OpenStudy (anonymous):

Rise means how much the graph goes up, or how much y increases. Run means how much the graph goes over, or how much x increases.

OpenStudy (anonymous):

Point slope form Y2-Y1 = M(X2-X1) 1+3 = M (-5-2) 4=M(-7) 4/-7=M A